### v3

@@ Abstract
[Abstract] Technical debt (TD) represents the long-term costs associated with suboptimal design or code decisions in software development, often made to meet short-term delivery goals. Self-Admitted Technical Debt (SATD) occurs when developers explicitly acknowledge these trade-offs in the codebase, typically through comments or annotations. Automated detection of SATD has become an increasingly important research area, particularly with the rise of natural language processing (NLP), machine learning (ML), and deep learning (DL) techniques that aim to streamline SATD detection.
This systematic literature review provides a comprehensive analysis of SATD detection approaches published between 2014 and 2024, focusing on the evolution of techniques from NLP-based models to more advanced ML, DL, and Transformers-based models such as BERT. The review identifies key trends in SATD detection methodologies and tools, evaluates the effectiveness of different approaches using metrics like precision, recall, and F1-score, and highlights the primary challenges in this domain, including dataset heterogeneity, model generalizability, and the explainability of models.
The findings suggest that while early NLP methods laid the foundation for SATD detection, more recent advancements in DL and Transformers models have significantly improved detection accuracy. However, challenges remain in scaling these models for broader industrial use. This SLR offers insights into current research gaps and provides directions for future work, aiming to improve the robustness and practicality of SATD detection tools.
@@ 1 Introduction
Business needs often require developers to be able to deliver new products and features as rapidly as possible, in order to reach the market. In fact, a developer must be able to balance carefully designing applications and seamlessly coding their features. In the business or corporate world, however, schedules and limited resources frequently hinder developers from generating perfect code before deploying it into products 1. This becomes commonplace when stakeholders cut several technical processes to catch up on momentum, especially in startup practices 2, 3. A tight deadline frequently leads to technical debt, as developers lack time to consider long-term solutions 4.
Technical debt (TD) is often used to describe the trade-off that occurs when an organization (or a developer) chooses to prioritize the delivery of new features or functionality over the long-term maintenance of the software 5. TD can occur in any software development project and can be caused by a variety of factors, including lack of resources, time, knowledge, or experience on the part of the development team 6, 7.
Among various types of TD, self-admitted technical debt (SATD) has gained considerable attention due to its explicit acknowledgment within the source code comments by developers. SATD represents instances where developers intentionally or unintentionally leave behind suboptimal design or implementation choices, often with a clear understanding of the resultant negative impact on the codebase 8. This acknowledgment is not merely an admission of suboptimal practices but a vital piece of information that can guide future development activities. Recognizing SATD is crucial for maintaining software quality and sustainability, as it can accumulate over time and impede future development efforts.
Over the past decade, many approaches to detecting SATD have emerged, driven by advances in natural language processing (NLP), machine learning (ML), and deep learning (DL). These approaches aim to automate the detection of SATD, facilitate their management, and helping software teams address it proactively. Alongside these approaches, various methodologies have been proposed to detect SATD in different stages of the software development life cycle, leveraging manual and semi-automated, and automated techniques. However, the diversity of techniques, methodologies, tools, and evaluation metrics used in SATD detection has led to a fragmented landscape, making it challenging for researchers and practitioners to gain a holistic understanding of the field.
@@
Our review is the most extensive to date, covering relevant studies published between 2014 and 2024. In contrast, the review by Sierra et al. focuses on a narrower time frame, covering studies published between 2014 and 2018 9.
@@
We followed the original and rigorous procedures outlined by Kitchenham and Charters 10 for conducting systematic reviews. In contrast, Sierra et al.’s study 9 appears to be based on a survey approach, without explicitly detailing how studies were sourced.
@@
In addition to relying on search engines, our review included a bidirectional snowballing process to ensure we captured all relevant articles. As a result, we analyzed a broader range of studies.
Profoundness.
Unlike Sierra et al. 9, who did not report the metric evaluation results, our review includes detailed analysis of metrics such as precision, recall, and F1-score, categorizing them based on the approaches and methodologies used.
Analysis.
While Sierra et al. 9 broadly discussed SATD activities without focusing deeply on detection approaches, our review provides a more detailed evaluation of detection techniques, their effectiveness, and the tools proposed for SATD detection.
In this study, we conduct and evaluate a systematic literature review that focuses on the detection of SATD, particularly using NLP-based approaches and their combinations. The objectives and key contributions of our study are as follows:
  • We consolidate and categorize various approaches employed to detect SATD, revealing various techniques and identifying critical trends over the past decade.
  • We examine methodologies proposed for SATD detection, showcasing their impact on automation and accuracy.
  • We investigate tools developed for SATD detection over the past decade, providing an overview of their effectiveness.
  • We compare the best-performing approaches by evaluating precision, recall, F1-score, and generalizability across different software systems and datasets.
  • We identify primary challenges in SATD detection, offering insights for future research and improvement.
In addition to the Introduction, the remaining study sections are organized as follows: Section 2 introduces the concept and background, and Section  3 describes the methodology utilized to conduct the SLR. Sections 4 and 5 summarize and discuss the obtained results, while Section 6 identifies the threats to validity of this study. Finally, Section 7 offers a conclusion.
@@ 2 Background
This section introduces the concepts underpinning this research, related to NLP and self-admitted technical debt, as well as the relevant works.
@@ 2.1 Self-Admitted Technical Debt (SATD)
The term “technical debt” was coined by software developer Ward Cunningham. He initially used the metaphor for communicating to non-technical stakeholders the importance of budgeting resources for refactoring 5. Technical debt, sometimes referred to as tech debt or code debt, is the result of a development team rushing the delivery of projects or features that must later be refactored. In other words, there is a trade-off between delivering high-quality, effective, and optimized code and the short-term benefits of achieving release dates 6, 4. A small amount of debt is acceptable because it can increase productivity, but if left uncontrolled, it could end up costing a lot of time, extra work, and money in the future 11, 12.
@@
The term “not-quite-right” is commonly used to refer to internal software development activities that are postponed but may lead to future difficulties if not addressed 6, 13. It represents the debt incurred when development teams opt for quick and easy solutions in the short term, risking long-term negative impacts.
The term “TD” encompasses below-par aspects of the software that developers are aware of but lack resources to address, such as outdated documentation, unperformed tests, complex code needing refactoring, and known defects left unattended 14, 15. These immature artifacts can lead to challenges in meeting quality standards and cause unforeseen delays in implementing necessary changes. Numerous empirical studies have recently concentrated on examining source code comments and other textual artifacts that indicate technical debt, often known as self-admitted technical debt (SATD) 8. The source used to detect SATD is usually in the form of natural language often expressed through text, such as code comments, commit messages, or other textual artifacts; therefore, SATD detection significantly relies on NLP-based approaches.
2.2 Natural Language Processing (NLP)
Natural Language Processing (NLP) is an interdisciplinary field that studies the interaction between computers and human language, encompassing tasks like speech recognition, natural language understanding, and natural language generation 16, 17. Chowdhury 17 explains that the objective of NLP researchers is to understand how people perceive and use language so that the appropriate tools and techniques can be developed to assist computers in understanding and using natural languages to complete the necessary tasks.
NLP deals with processing natural language, which is the language used by humans in text or audio form. Computational linguistics, rule-based human language modeling, statistical models, machine learning, and deep learning are combined to enable computer programs to interpret and process human language 18, 19. In the context of SATD detection, plain text is used as input, including project artifacts such as source code comments, commit messages, pull requests, and issue tracking systems 20, 21.
Source code comments serve as annotations within the code, providing documentation directly incorporated by programmers 22. These comments are essential for understanding a system and may contain important information, such as the purpose of code additions, progress status on collective tasks, or reasons for specific changes 23. Comments can express concerns that need attention in the future (e.g., the TODO keyword), identify emerging issues, and guide future actions 24, 25. They provide valuable context specific to their authors and are human-readable.
Recently, automated approaches utilizing NLP have gained traction in addressing software engineering problems, including the detection of TD, which is called SATD detection. NLP has been effectively used to automatically identify the presence of TD and even categorize specific types of SATD. Notably, numerous studies have integrated NLP with machine learning (ML) models or deep learning (DL) architectures and also Transformers, aiming to enhance model performance (evidence is provided in Subsection 4.1).
@@ 3 Research Methodology
This study is carried out in accordance with the methodology of conducting an SLR described by Kitchenham et al26, 27, 10. The procedures for SLR are broken down into three stages: planning, conducting, and reporting the SLR 10. We developed a review protocol during the planning stage that included the following steps: i) identification of research topics, ii) creation of a search strategy, iii) study selection criteria, iv) study quality assessment, v) data extraction procedure, and vi) data synthesis process (as illustrated in Figure  1).
Figure 1: Graphical representation of the SLR process used in this article
The review process comprises several steps. Initially, research questions were formulated for the SLR. Subsequently, a search strategy was devised, involving the identification of search strings and the selection of online databases for locating primary studies. The third stage encompassed the identification of pertinent studies aligned with the research questions, along with the establishment of inclusion and exclusion criteria. Following this, bidirectional snowballing was conducted utilizing references from the source articles. After that, we created a data extraction form to answer the research questions. A data extraction form was then crafted to address the research questions. Lastly, a data synthesis approach was developed to analyze and summarize the research findings.
Figure 2 depicts the article selection procedure for determining the primary studies used in this SLR. The following subsections provide detailed information about the selection process.
Figure 2: An overview of the steps in determining primary studies
@@ 3.1 Research Questions
Our SLR includes all relevant research outlining the various SATD detection techniques. To define the search string, we systematically broke down the research questions using the PICOC framework (Population, Intervention, Comparison, Outcome, and Context), as recommended by Petticrew and Roberts 28. Boolean operators were applied to structure the search effectively: OR was used to include synonyms and alternative terms, while AND linked the key terms, as shown in Table 1. This approach ensured a comprehensive and focused search strategy.
Table 1: PICOC Criteria
@@
Comparison Not aplicable Not aplicable
@@
According to the PICOC in Table 1, the main objective of this SLR is to investigate, identify, and classify all relevant evidence from the primary studies that focus on using NLP to detect self-admitted technical debt. Therefore, the aim of this SLR is to answer the following questions:
RQ1: What approaches have been employed to detect self-admitted technical debt (SATD) in software development?
Rationale: This question aims to understand the progress made in detecting SATD. By surveying existing studies, we can detect trends and potential future research directions.
RQ2: What methodologies have been proposed to detect SATD in software development?
Rationale: This question aims to analyze the various approaches and tools developed over the past decade for SATD detection.
RQ3: How effective are these approaches regarding their metrics evaluation results?
Rationale: This question aims to evaluate the performance and practicality of the identified approaches by comparing their precision, recall, and F1-score.
@@ 3.2 Search Protocol
Before beginning the SLR planning stage, it is necessary to conduct a pilot search on a database such as Scopus to observe if there are any SLRs with the same research theme. There is no need to conduct a new SLR if there are existing SLRs on the same topic 29. In this systematic review of the literature, no results covering the topic of the capabilities of the NLP approach in identifying technical debt were discovered; therefore, the SLR can be conducted as new research.
The SLR’s search strategy consists of three stages: pilot search, primary search, and snowballing. In the pilot search, multiple search strings are used on online databases to determine the most effective one. The primary search uses the identified search string to find relevant literature in online source databases. The search concludes with a snowball search 30. The pilot and primary search utilized six online databases: Web of Science, Scopus, IEEE Xplore, ScienceDirect, ACM Digital Library, and Springer. These databases, as proved by previous software engineering SLRs 31, 32, 33, 34, provide an excellent overview of the software engineering literature and are trustworthy enough to cover our scope of the literature. Moreover, we employed the “Advanced Search” feature in these databases to obtain full-text and metadata results.
Table 2: General search strings utilized during pilot search
@@
S1 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”)
S2 (“self-admitted technical debt" OR SATD) AND (NLP OR “natural language processing”) AND (detection OR identification)
S3 (“self-admitted technical debt" OR SATD) OR (“technical debt" AND NLP) AND (detect* OR identif* OR predict*) AND (“software engineering" OR “software development")
The results of a pilot search using search string S1 were communicated to the second author by the first author. The second author selected 5% articles at random to see if they were relevant. The second author proposed altering the search string to S2 because the search string S1 result was still too general. Similar to the first process, iteration is repeated until it is decided to utilize the final search string, namely search string S3.
As illustrated in Table 2, through a pilot search with multiple search strings, the most effective one (i.e., S3) was selected. The search string was designed with synonyms, alternative keywords, and boolean operators ‘OR’ and ‘AND’ to retrieve relevant studies. Drawing from the pilot search conducted earlier, for this search string S3, we incorporated three supplementary terms commonly associated with tasks related to SATD identification: detect, identify, and predict. Additionally, wildcard asterisk (*) was used to broaden the search for variations of words 35.
The final search string (i.e., S3) was applied to the mentioned 6 online databases to find relevant primary studies for the research questions as described in Subsection 3.1. We utilized the “Advanced Search” feature in these databases to acquire full-text and metadata results.
Table 3: Search results from selected databases
@@
1 Web of Science 51
2 Scopus 1,004
3 IEEE Xplore 168
4 ACM Digital Library 41
5 Springer 125
6 ScienceDirect 51
Total 1,440
As depicted in Figure 2, the initial step involved searching multiple online databases using predefined search queries to collect a comprehensive set of studies relevant to the research topic. As shown in Table 3), a total of 1,440 studies were retrieved at this stage. To manage such a large dataset, both authors worked collaboratively to remove duplicates, which resulted in 1,119 unique records. Following the duplicate removal, a filtering process based on metadata by using inclusion and exclusion criteria presented in Table 4 was conducted. At this stage, the titles, abstracts, and keywords of the remaining studies were examined to determine their relevance to SATD detection.
A full-text analysis of these 69 studies was then performed to evaluate the quality and relevance of the research in greater detail. The first author initially reviewed the metadata and conducted an independent assessment of the full papers, focusing on their contributions to SATD detection approaches. After completing the initial evaluation, the first author discussed the findings with the second author, who served as a supervisor throughout the process. This discussion provided a critical review of the selection criteria and the individual assessments.
To ensure that no bias influenced the selection process, the second author independently reviewed a subset of the full-text studies. These results were then compared with the first author’s evaluations, and any discrepancies or differences in judgment were carefully discussed and resolved. This collaborative approach minimized the risk of subjective bias and ensured that each study was rigorously and objectively evaluated based on its quality and relevance.
To further ensure consistency in the evaluation process, the test-retest strategy outlined by Kitchenham and Charters 10 was employed, rather than relying on inter-rater agreement. This strategy involved discussions between the first and second authors to maintain a consistent evaluation framework. Additionally, the second author cross-checked the data extraction results, which were shared in a spreadsheet [^1] The spreadsheet is available in the shared dataset area at https://github.com/edisutoyo/satd-detection-slr. Any disagreements that arose during the process were addressed through collaborative discussions until consensus was reached.
After this comprehensive review and resolution of all discrepancies, both authors agreed to select 63 relevant studies from 6 online databases to be included in the systematic review. This thorough, unbiased, and transparent process ensured the inclusion of only the most relevant and high-quality studies.
3.3 Article Selection and Snowballing Process
Inclusion and exclusion criteria must be precisely defined for study selection. The inclusion criteria should include an automated approach to detect technical debt. Consequently, the primary studies were chosen based on comparing titles, keywords, abstracts, and full texts (i.e., introduction, methodology, results, and conclusion). We employed 6 inclusion criteria and 5 exclusion criteria, as shown in Table 4):
Table 4: Inclusion and exclusion criteria

### v4

@@ Abstract
[Abstract] Technical debt (TD) refers to the long-term costs associated with suboptimal design or code decisions in software development, often made to meet short-term delivery goals. Self-Admitted Technical Debt (SATD) occurs when developers explicitly acknowledge these trade-offs in the codebase, typically through comments or annotations. SATD detection has become an increasingly important research area, particularly with the rise of learning-based techniques that aim to streamline SATD detection.
This systematic literature review provides a comprehensive analysis of SATD detection approaches published between 2014 and early 2025, focusing on the evolution of techniques from heuristic-based techniques to more advanced ML, DL, and Transformer-based models. It examines key trends in SATD detection methodologies and tools, evaluates the effectiveness of different approaches using metrics like precision, recall, and F1 score, and highlights the primary challenges in this domain, including dataset heterogeneity, model generalizability, and explainability.
The findings reveal that while early heuristic-based techniques laid the foundation for SATD detection, more recent advancements in DL and Transformer models have significantly improved detection accuracy. However, challenges remain in scaling these models for broader industrial adoption. This review offers insights into current research gaps and provides directions for future work, aiming to improve the robustness and practicality of SATD detection tools.
@@ 1 Introduction
Business needs often require developers to rapidly deliver new products and features to reach the market quickly. It is important for developers to carefully balance design and implementation to seamlessly add features. However, in the business or corporate world, schedules and limited resources often hinder developers from writing perfect code before incorporating it into products 1. This becomes commonplace when stakeholders cut technical processes to gain momentum, especially in startups 2, 3. Tight deadlines often lead to technical debt because developers lack time to consider long-term solutions 4.
Technical debt (TD) is often used to describe the trade-off that occurs when an organization prioritizes delivering new features or functionality over maintaining the software in the long term 5. TD can occur in any software development project and may be caused by a variety of factors, including insufficient resources, time, knowledge, or experience on the part of the development team 6, 7.
Among various types of TD, self-admitted technical debt (SATD) has gained considerable attention due to its explicit acknowledgment within the source code comments by developers. SATD refers to instances in which developers, either intentionally or unintentionally, make suboptimal design or implementation choices, despite having a clear understanding of the negative impact these choices will have on the codebase 8. This acknowledgment is not merely an admission of suboptimal practices but a vital piece of information that can guide future development activities. Recognizing SATD is crucial for maintaining software quality and sustainability, as it can accumulate over time and impede future development efforts.
Over the past decade, various approaches to SATD detection have emerged, evolving from heuristic-based techniques to more advanced machine learning (ML), deep learning (DL), and Transformer-based models. These approaches aim to automate SATD detection, facilitate its management, and enable software teams to address it proactively. Alongside these approaches, various methodologies have been proposed to detect SATD, ranging from manual to semi-automated to automated methodologies. However, the diversity of techniques, methodologies, tools, and evaluation metrics used in SATD detection has led to a fragmented landscape, making it challenging for researchers and practitioners to gain a holistic understanding of the field.
@@
Our review spans a decade-long period from January 2014 to April 2025, analyzes 74 studies on SATD detection, and provides the most up-to-date synthesis of SATD research. In contrast, the review by Sierra et al. 9 focuses on a shorter interval (2014–2018) and captures earlier SATD detection efforts. Similarly, Gu et al. 10 conduct a lightweight review that covers literature until 2022, but primarily focuses on evaluating detection models across different software artifacts rather than providing a broad literature synthesis.
@@
We follow the established and rigorous procedures outlined by Kitchenham and Charters 11 for conducting systematic reviews. In contrast, Sierra et al. 9 appears to be based on a survey approach, without explicitly detailing how studies are sourced. Gu et al. 10 conduct an empirical study, but their review methodology does not follow an SLR framework, as it primarily assesses the generalization of SATD models across programming languages and software artifacts.
@@
In addition to relying on digital library searches, our review employs a bidirectional snowballing process to capture all relevant studies and results, resulting in 74 analyzed papers and making our review broader and more representative of the field. While Sierra et al. 9 also employ snowballing, their review covers a narrower time frame (2014–2018), which may limit the discovery of additional relevant studies. In contrast, Gu et al. 10 do not explicitly perform snowballing, primarily focusing on evaluating existing SATD detection models rather than conducting a comprehensive literature search.
Depth of analysis.
Prior reviews, such as Sierra et al. 9 and Gu et al. 10, do not consistently synthesize reported findings. Our review extends this by examining both methodological characteristics (e.g., degree of automation) and empirical evidence (e.g., reported performance metrics), offering a more integrated perspective on how SATD detection techniques evolve and perform.
In this study, we conduct a systematic literature review on SATD detection, evaluating existing approaches and their effectiveness. The objectives and key contributions of our study are as follows:
  • We consolidate and classify SATD detection approaches, tracing the evolution from heuristic-based methods to ML, DL, and Transformer-based models. This taxonomy provides researchers with a structured view of methodological progress in the field.
  • We compare methodologies in terms of their level of automation (manual, semi-automated, and automated) and synthesize reported performance metrics (precision, recall, and F1 score) across diverse datasets and software systems. This synthesis provides a consolidated overview of the techniques and highlights the best-performing models reported in prior studies.
  • We conduct the first systematic review of SATD detection tools, assessing their availability, capabilities, and reported effectiveness. This analysis reveals persistent gaps between research prototypes and practical tool adoption in industry settings.
  • We identify and synthesize nine recurring challenges in SATD detection (e.g., dataset imbalance, model generalizability and explainability, and adoption barriers), which highlight key opportunities for future research.
In addition to the Introduction, the remaining study sections are organized as follows: Section 2 introduces the concept and background, and Section 3 describes the methodology utilized to conduct the SLR. Sections 4 and 5 summarize and discuss the obtained results, while Section 6 identifies the threats to validity of this study. Finally, Section 7 offers a conclusion.
@@ 2 Background
This section introduces the key concepts underlying this study, focusing on self-admitted technical debt (SATD) and the role of natural language processing (NLP) techniques that are directly relevant for analyzing the textual artifacts used in SATD detection.
@@ 2.1 Self-Admitted Technical Debt (SATD)
The term “technical debt” (TD) was coined by software developer Ward Cunningham. He initially used the metaphor for communicating to non-technical stakeholders the importance of budgeting resources for refactoring 5. Technical debt, sometimes referred to as tech debt or code debt, is the result of a development team rushing the delivery of projects or features that must later be refactored. In other words, there is a trade-off between delivering high-quality, effective, and optimized code and the short-term benefits of achieving release dates 6, 4. A small amount of debt is acceptable because it can increase productivity, but if left uncontrolled, it could end up costing a lot of time, extra work, and money in the future 12, 13. The “debt” metaphor further implies that, just as financial debt accrues interest, the longer suboptimal code remains in the system, the more costly and time-consuming it becomes to remove. This emphasizes the compounding nature of technical debt over time, where delayed repayment leads to greater maintenance burdens and reduced productivity.
@@
The concept “not-quite-right” is commonly used to refer to internal software development activities that are postponed but may lead to future difficulties if not addressed 6, 14. It represents the debt incurred when development teams opt for quick and easy solutions in the short term, risking long-term negative impacts.
TD encompasses various suboptimal aspects of software that developers acknowledge but lack the immediate resources to address. These may include outdated documentation, missing tests, complex code in need of refactoring, and known defects left unresolved 15, 16. Such deficiencies can degrade software quality over time, making maintenance more challenging and causing unforeseen delays when implementing necessary changes.
In recent years, numerous empirical studies have focused on identifying technical debt through textual artifacts, particularly source code comments and other developer-written text, a practice commonly referred to as self-admitted technical debt (SATD) 8. The primary sources for SATD detection include code comments, commit messages, issue tracker discussions, and other natural language-based software artifacts.
As SATD is expressed in natural language, its detection relies heavily on language processing techniques to analyze and classify technical debt instances. Therefore, SATD detection can be regarded as an application of NLP within the broader context of software engineering, as it relies heavily on natural language techniques to identify and classify SATD.
2.2 NLP in SATD Detection
Natural Language Processing (NLP) is an interdisciplinary field that studies the interaction between computers and human language. It includes tasks such as speech recognition, natural language understanding, and natural language generation 17, 18. Within the scope of SATD detection, NLP is applied in a more focused way as it provides the means to transform developer-authored textual artifacts such as code comments, commit messages, issue tracker discussions, and pull requests into structured features that automated models can analyze 19, 20. For instance, code comments often contain explicit markers of pending work (e.g., TODO) or implicit expressions of design compromise, which can be mined using NLP-based approaches 21, 22, 23.
Unlike traditional static code analysis approaches that detect proxies of technical debt, such as code smells or rule violations, SATD detection aims to uncover technical debt as explicitly acknowledged by developers themselves 8, 24. This makes SATD a unique and valuable complement to other TD detection methods, as it directly reflects developers’ awareness of design trade-offs and postponed tasks 25.
In prior studies, SATD detection has been approached from two complementary perspectives. Some research focuses on determining whether a given software artifact contains any indication of debt 24, 26, thereby providing an overview of its extent within a system. Other research takes this further by distinguishing between different forms of debt, such as design, architecture, requirement, or test debt, based on the linguistic patterns found in the text 24, 27, 28. This finer-grained view is particularly useful because different types of SATD have distinct implications for maintenance and repayment strategies 9. Together, these perspectives highlight how detection approaches can support both the recognition and the deeper understanding of technical debt in practice 29.
Over time, SATD detection techniques have evolved in line with advances in NLP, moving from simple keyword-based heuristics to classical machine learning classifiers trained on TF-IDF features, and more recently to deep learning and Transformer-based models that leverage semantic embeddings 24, 30, 31. Although these approaches demonstrate significant progress, the field is still maturing and continues to face open questions that require systematic investigation. This underlines the importance of consolidating and critically examining the state of the art, which is the focus of this review.
While the background section introduced SATD and its detection techniques, a systematic investigation is required to consolidate existing research findings. To achieve this, we employ a structured literature review methodology. The following section outlines our research design, including our research questions, study selection process, and data extraction methods, ensuring a rigorous and reproducible review.
@@ 3 Research Methodology
This study is carried out in accordance with the methodology of conducting an SLR described by Kitchenham et al. 32, 33, 11. The procedures for SLR are divided into three stages: planning, conducting, and reporting 11. We developed a review protocol during the planning stage that included the following steps: i) identification of research topics, ii) creation of a search strategy, iii) study selection criteria, iv) study quality assessment, v) data extraction procedure, and vi) data synthesis process (as illustrated in Figure 1).
Figure 1: Graphical representation of the SLR process used in this article.
The review process comprises several steps. Initially, research questions were formulated for the SLR. Subsequently, a search strategy was devised, involving the development of search strings and the selection of online databases to locate primary studies. The third stage encompassed the identification of pertinent studies aligned with the research questions, along with the establishment of inclusion and exclusion criteria. To ensure comprehensive coverage, a bidirectional snowballing approach was employed. This involved examining the references cited in the source articles and exploring articles that cited these sources, using Google Scholar [^1] https://scholar.google.com as the primary tool, as recommended in prior guidelines 34. Subsequently, we developed a data extraction form to address the research questions. A data extraction form was then crafted to address the research questions. Lastly, a data synthesis approach was developed to analyze and summarize the research findings.
@@ 3.1 Research Questions
Our SLR reviews all relevant research outlining the various SATD detection techniques. To define the search string, we systematically broke down the research questions using the PICOC framework (Population, Intervention, Comparison, Outcome, and Context), as recommended by Petticrew and Roberts 35. Boolean operators were applied to structure the search effectively: OR was used to include synonyms and alternative terms, while AND linked the key terms, as shown in Table 1. This approach ensured a comprehensive and focused search strategy.
Table 1: PICOC Criteria.
@@
Comparison Not applicable Not applicable
@@
According to the PICOC in Table 1, the main objective of this SLR is to investigate, identify, and classify all relevant evidence from the primary studies that focus on using NLP to detect SATD. Therefore, this SLR aims to answer the following questions:
RQ1: What types of techniques have been employed to detect SATD in software development, and how do these techniques handle different types of textual artifacts?
Rationale: This question aims to categorize and analyze the various types of techniques that have been proposed for SATD detection across different types of textual artifacts, providing insight into trends and gaps in the field.
RQ2: What levels of automation have been proposed in methodologies for detecting SATD in software development, and to what extent do these methods rely on human interventions across different types of textual artifacts?
Rationale: This question focuses on understanding the extent to which SATD detection methods incorporate automation, highlighting the progress and identifying gaps in automating the detection process across different types of textual artifacts.
RQ3: Which SATD detection techniques are reported as the most effective based on evaluation metrics?
Rationale: This question aims to review and synthesize the reported precision, recall, and F1 scores of the identified techniques to provide an overview of their observed performance. The goal is not to perform an independent benchmark but to consolidate how techniques have been reported in the literature and to identify which are most frequently reported as achieving the highest effectiveness.
@@ 3.2 Search Protocol
Before initiating the SLR planning stage, it is essential to conduct a pilot search in databases such as Scopus [^2] https://www.scopus.com to determine whether existing SLRs have already addressed the same research theme. If a relevant SLR is found, conducting a new one may not be necessary 36.
Our preliminary search revealed no existing SLRs encompassing the breadth of approaches for detecting SATD. Consequently, this SLR fills a critical gap in the literature and offers a timely contribution to the field.
The SLR’s search strategy consists of three stages: pilot search, primary search, and snowballing. In the pilot search, multiple search strings are used in online databases to identify the most effective. The primary search uses the identified search string to retrieve relevant literature from online databases. The search concludes with a snowball search 37. The pilot and primary search utilized six online databases: Web of Science, Scopus, IEEE Xplore Digital Library, ScienceDirect, ACM Digital Library, and Springer. These databases, as demonstrated by previous software engineering SLRs 38, 39, 40, 41, provide an excellent overview of the software engineering literature and are sufficiently comprehensive to cover our scope. Moreover, we employed the “Advanced Search” feature in these databases to obtain full-text and metadata results.
The first author communicated the results of a pilot search using the search string S1 to the second author. The second author randomly selected 10% of the articles to check their relevance. Based on the results, the second author proposed modifying the search string to S2 because the results from S1 remained too general. Similar to the initial process, the iteration was repeated until it was decided to use the final search string, namely S3 (see Table 2).
Table 2: General search strings utilized during pilot search.
@@
S1 (“self-admitted technical debt” OR SATD) AND (NLP OR “natural language processing”)
S2 (“self-admitted technical debt” OR SATD) AND (NLP OR “natural language processing”) AND (detection OR identification)
S3 (“self-admitted technical debt” OR SATD) OR (“technical debt” AND NLP) AND (detect* OR identif* OR predict*) AND (“software engineering” OR “software development”)
Through a series of pilot searches with multiple search strings, the most effective one (i.e., S3) was selected. The search string was designed with synonyms, alternative keywords, and the Boolean operators ‘OR’ and ‘AND’ to retrieve relevant studies. Building on the earlier pilot search, we expanded search string S3 by adding three supplementary terms that are commonly used in SATD identification tasks: detect, identify, and predict. Additionally, the wildcard asterisk (*) was used to broaden the search for variations of words 42.
The final search string (i.e., S3) was then applied to the six selected online databases to identify relevant primary studies addressing the research questions described in Subsection 3.1. We utilized the “Advanced Search” feature in these databases to obtain full-text and metadata results, as summarized in Table 3.
Table 3: Search results from selected databases.
@@
1 Web of Science 56
2 Scopus 1,118
3 IEEE Xplore 184
4 ACM Digital Library 49
5 Springer 181
6 ScienceDirect 69
Total 1,659
3.3 Study Selection
As shown in Figure 2, we selected primary studies through a staged screening procedure to retain only relevant and sufficiently detailed evidence for this review. The initial search across multiple digital libraries returned 1,659 records. After duplicate removal, 1,245 unique records remained. We then applied a three-stage screening process guided by the inclusion and exclusion criteria in Table 4.
Figure 2: An overview of the steps in determining primary studies.
First, we performed a metadata screening of titles, abstracts, and keywords to assess topical relevance, which reduced the set to 194 studies. Second, we screened studies for research methodology and eligibility (e.g., whether the study reported an SATD identification approach and provided sufficient methodological detail), yielding 85 studies. Third, we conducted full-text screening to confirm relevance and assess the study’s contribution and reporting quality, yielding 69 studies focused on SATD detection. These 69 studies constitute the initial core set of primary studies for our SLR.
Table 4: Inclusion and exclusion criteria.